Find the 90th. Percentile of the t-distribution if the sample size is 25?







1
Expert's answer
2022-06-10T14:17:51-0400

Based on the information provided, the significance level is "\alpha = 0.10," "df=n-1=24" degrees of freedom.

The 90th Percentile "t=1.710882."